OpenStudy (anonymous):

Matthew is waterskiing. As the boat starts moving, he is at an angle of 8.0° to the right of the boat. The boat applies 250 newtons over the first 50 meters, while the angle stays at roughly 8.0°. What work has the boat done on Matthew?

OpenStudy (mayankdevnani):

workdone=Fd*cos(theta) =250*50*cos(8) =12500*cos(8)=12378.35joule ok @cbowser1211

OpenStudy (mayankdevnani):

got it

OpenStudy (anonymous):

thanks for the help are you on here all day lol

OpenStudy (mayankdevnani):

yaa